Last q before my exam: Football fans claim that it takes 35 minutes to get out of the stadium after a game when over 30000 fans attend the game.
To determine if the complains are justified , a random sample of 100 fans is taken. The average time it takes these 100 fans to exit the parking is 27 minutes and sample standard deviatian is 7 minutes.
a) Estimate the mean departe time= 27 minutes?
b) Construct a 95% CI for the mean departure time and interpret the result?

Thx soo much for all your help

Answer by stanbon(48568)

Football fans claim that it takes 35 minutes to get out of the stadium after a game when over 30000 fans attend the game.
To determine if the complains are justified , a random sample of 100 fans is taken. The average time it takes these 100 fans to exit the parking is 27 minutes and sample standard deviatian is 7 minutes.
a) Estimate the mean departe time = 27 minutes?
b) Construct a 95% CI for the mean departure time and interpret the result?
----------
sample mean: 27 min
------
standard error: SE = t*s/sqrt(n)
t = invT(0.025,99) = 1.9842
SE = 1.9842*7/sqrt(100) = 1.389
---
95% CI: 27-1.389 < u < 27+1.389
95% CI: 25.61 < u < 28.389
---------------------------------
Meaning: We are 95% confident the mean departure time is between
25.61 min and 28.389 min
==================================
Cheers,
Stan H.